The partnership of Anderson, Berry, Hammond, and Winwood is being liquidated. It currently holds cash of $22,500 but no other assets. Liabilities amount to $35,600. The capital balances are

Anderson (40% of profits and losses) $ 27,900
Berry (20%) 13,000
Hammond (20%) (25,000 ) (deficit)
Winwood (20%) (29,000 ) (deficit)
  1. a. If both Hammond and Winwood are personally insolvent, how much money must Berry contribute to this partnership?

  2. b-1. If only Winwood is personally insolvent, how much money must Hammond contribute?

  3. b-2. If only Winwood is personally insolvent, how will Hammond's contribution (funds) be disbursed?

  4. c. If only Hammond is personally insolvent, how much money should Anderson receive from the liquidation?
Answer #1

Answer a.

Berry contribute to this partnership = [(25000+29000)*20/60] - 13000 = 5000

Answer b1.

Hammond contribute = [ 29000 * 20/80 ] +25000 = 32250

Answer b2.

Creditors = 13100

Anderson = [29000 *40/80] - 27900 = 13400

Berry = [29000 *20/80] - 13000 = 5750

Answer c.

= 25000 *40/80 = 12500

= 27900 -12500

= 15400